The unadjusted trial balance of Peace Hill Inc. at December 31, 2017, is as follows:

Debit Credit $17,120 103,100 Cash Accounts Receivable Allowance for Doubtful Accounts Inventory Prepaid Insurancee Bond Investment at Amortized Cost Land Buildings Accumulated Depreciation-Buildings Equipment Accumulated Depreciation-Equipment Goodwill Accounts Payable Bonds Payable (20-year, 8%) Common Shares Retained Earnings Sales Revenue Rent Revenue Advertising Expense Supplies Expense Purchases Purchase Discounts Salaries and wages expense Interest Expense $3,490 60,400 4,636 57,120 31,700 155,900 6,205 33,240 5,540 16,700 100,200 174,000 115,100 68,976 187,500 10,425 20,250 10,200 97,000 930 53,000 12,000 $672,366 $672,366

Additional information:

1. Actual advertising costs amounted to $1,350 per month. The company has already paid for advertisements in Crizza MMagazine for the first quarter of 2018.
2. The building was purchased and occupied on January 1, 2015, with an estimated useful life of 20 years, and residual value of $31,800. (The company uses straight-line depreciation.)
3. Prepaid insurance contains the premium costs of several policies, including Policy A, cost of $2,568, one-year term, taken out on April 1, 2017; and Policy B, cost of $2,068, three-year term, taken out on September 1, 2017.
4. A portion of Peace Hill’s building has been converted into a snack bar that has been rented to the ABC Corp. since July 1, 2016, at a rate of $6,950 per year payable each July 1.
5. One of the company’s customers declared bankruptcy on December 30, 2017. It is now certain that the $2,760 the customer owes will never be collected. This fact has not been recorded. In addition, Peace Hill estimates that 5% of the Accounts Receivable balance on December 31, 2017, will become uncollectible.
6. An advance of $620 to a salesperson on December 31, 2017, was charged to Salaries and Wages Expense.
7. On November 1, 2015, Peace Hill issued 174 $1,000 bonds at par value. Interest is paid semi-annually on April 30 and October 31.
8. The equipment was purchased on January 1, 2015, with an estimated useful life of 10 years, and no residual value. (The company uses straight-line depreciation.)
9. On August 1, 2017, Peace Hill purchased at par value 42 $1,360, 8% bonds maturing on July 31, 2019. Interest is paid on July 31 and January 31.
10.

The inventory on hand at December 31, 2017, was $88,800 after a physical inventory count. (Use "Inventory" account for closing out the beginning inventory amount and recording the ending inventory amount.)

Prepare adjusting and correcting entries for December 31, 2017, using the information given. Record the adjusting entry for inventory using a Cost of Goods Sold account.

No. Account Titles and Explanation Debit Credit 2. 3. 4. 5. To write off uncollectible receivables.) (To record bad debt expense.)

6. 7. 8. 9. 10

HERE ARE THE CHOICE OF ACCOUNT TITLES TO USE:

Accounts Payable
Accounts Receivable
Accrued Liabilities
Accumulated Amortization - Patent
Accumulated Depreciation - Buildings
Accumulated Depreciation - Equipment
Accumulated Depreciation - Furniture and Fixtures
Accumulated Depreciation - Machinery
Accumulated Depreciation - Trucks
Accumulated Impairment Losses
Accumulated OCI
Administrative Expenses
Advances to Employees
Advertising Expense
Allowance for Doubtful Accounts
Amortization Expense
Bad Debt Expense
Buildings
Cash
Common Shares
Cost of Goods Sold
Current Portion of Notes Payable
Deferred Development Costs
Delivery Expense
Depreciation Expense
Dividends
Dividends Payable
Equipment
Fair Value - NI Investments
Fair Value - OCI Investments
Freight-in
Freight-out
Furniture and Fixtures
Impairment Loss
Income Summary
Income Tax Expense
Insurance Expense
Insurance Revenue
Interest Expense
Interest Income
Interest Payable
Interest Receivable
Inventory
Investment Income or Loss
Land
Legal Expense
Loss on Inventory Due to Decline in NRV
Miscellaneous Expense
Mortgage Payable
No Entry
Notes Payable
Notes Receivable
Office Expense
Operating Expenses
Owner's Capital
Owner's Drawings
Patents
Prepaid Advertising
Prepaid Expenses
Prepaid Insurance
Prepaid Rent
Property Tax Expense
Property Tax Payable
Purchase Discounts
Purchase Returns and Allowances
Purchases
Rent Expense
Rent Payable
Rent Receivable
Rent Revenue
Repairs and Maintenance Expense
Repair Revenue
Retained Earnings
Salaries and Wages Expense
Salaries and Wages Payable
Sales Commission Expense
Sales Commission Payable
Sales Discounts
Sales Returns and Allowances
Sales Revenue
Service Revenue
Supplies
Supplies Expense
Telephone Expense
Trucks
Unearned Rent Revenue
Unearned Repair Revenue
Unearned Revenue
Unrealized Gain or Loss - OCI
Utilities Expense
Utilities Payable
Answer #1
No. A/c Titles & Explanations Debit Credit
1 Prepaid Advertising 4050
Advertising expense 4050
(1350*3)
2 Retained earnings (for 2016 depn.) 6205
Depreciation Expense 6205
Accumulated Depreciation-Buildings 12410
(155900-31800)/20
3 Insurance expense 2156
Prepaid Insurance 2156
Policy A-(2568/12*9)+
Policy B- (2068/36*4)
4 Rent Revenue 3475
Unearned Rent Revenue 3475
6950/12*6 mths.
5 Allowance for Doubtful Accounts 2760
Accounts Receivable 2760
(to record Uncollectible receivables)
Bad debt expense 4287
Allowance for Doubtful Accounts 4287
(103100-2760)=100340*5%=5017+2760-3490
(To record additional bad debt expense provision required)
6 Advances to Employees 620
Salaries and wages expense 620
7 Interest expense 1920
Interest payable 1920
(174000*8%)-12000
8 Retained earnings (Prior Year) 1108
Depreciation expense 3324
Accumulated Depreciation-Equipment 4432
(33240/10)*3 yrs.=9972-5540=4432
9 Interest receivable 1904
Interest Income 1904
(57120*8%*5 mths./12)
10 Inventory(Closing) 88800
COGS(Bal.fig.) 67670
Purchase Discounts 930
Purchases 97000
Inventory(Beginning) 60400